OpenStudy (maxdrizner):

What is the midpoint of the line segment with endpoints (5, -2) and (1, -6)? (2, -2) (3, -4) (3, -6) (4, -3)

hartnn (hartnn):

The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (maxdrizner):

thank you

hartnn (hartnn):

welcome ^_^